OpenStudy (anonymous):

Verify that y1(t)=t^2 and y2(t)=t^-1 are two solutions of the differential equation t^2y''-2y=0 for t>0. Then show that y=C1t^2+C2t^-1 is also a solution of this equation for any C1 and C2

OpenStudy (jamesj):

So substitute in and show it works.

OpenStudy (anonymous):

I don't understand.. What is it that I'm supposed to substitute in?

OpenStudy (jamesj):

You want to show that y1 and y2 satisfy the equation t^2y''-2y=0 So, for y1, for instance, y1 = t^2, y1' = 2t, y1'' = 2 thus t^2y''-2y = t^2 . 2 - 2 . t^2 = 0 hence y1 is a solution.

OpenStudy (jamesj):

Now do the same thing with y2 ... ... and then with the sum of those terms.

OpenStudy (anonymous):

THANKS SO MUCH!

OpenStudy (anonymous):

what about the second part?